A double acting steam engine runs at 210 rpm. A curve of turning moment plotted on the crank angle base showed the following areas alternatively above and below the mean turning moment line: 880,500,620,720,360,560,440,520mm2

The scales used where 1mm= 520Nm and 1mm= 1o crank angle. If the total fluctuation in speed is limited to 1.85% of the ,mean speed, determine the mass of flywheel necessary if the radius gyration is 1.4m



1
Expert's answer
2021-03-11T06:57:00-0500

Maximum energy ="880+E", where E is the minimum energy

"\triangle E=E_{max}-E_{min}=E +880-E=880 mm^2"

"\triangle E=880 \times520 \times \frac{\pi}{180}=7986.62666 Nm"

We know that "\triangle E =mR^2w^2C_s"

"w_1=210 \times1.85 \times \frac{2 \pi}{60}=40.68"

"w_2=210 \times \frac{2 \pi}{60}=10.995"

"w=\frac{w_1+w_2}{2}=\frac{40.68+10.995}{2}=25.8375"

"C_s=\frac{w_1-w_2}{w}=\frac{40.68-10.995}{25.8375}=1.1489"

"\triangle E =mR^2w^2(\frac{w_1-w_2}{w})=mR^2w(w_1-w_2)"

"m=\frac{\triangle E}{R^2w(w_1-w_2)}"

"m=\frac{7986.62666}{1.4^2 \times 25.8375(40.68-10.995)}=5.313 kg"